please show all steps and label them and also calculate the b values(1 point) The second order equation 3xy- 4xy + (x2 + 2)y = 0 has a regular singular point at x = 0, and therefore has a serie
Answer #1

+ E 2₂ wnan Cnanta in series holn neo ись Let 919)x? Ź inan Ž Cnlntv) (-1) 3x² Y-4xy +12+2) Y=0 (1) x=0 in regular singularwe get Co - 6+2)(36-1)+2 using (t) for m=2, 3, 4, C2 = (v +2) (394 2 +38-7) +2 C4 = C2 (v+4)(31 3r+5) +2 (37+5)+23 -(0) Со C6For any doubt please mention in comment box. Thanks